Zeronknight (zeronknight):

If Becky can fold 12 sweaters in 15 minutes, how long will it take her to fold 20 sweaters? * 25 minutes 20 minutes 15 minutes 18 minutes

OpenStudy (legomyego180):

\[\frac{ 12 }{ 15 } = \frac{ x }{ 20 }\]

OpenStudy (legomyego180):

(12*20)/15= sweaters she can fold in 20 min. This is a proportion.

Zeronknight (zeronknight):

12 x 20 = 240?

OpenStudy (legomyego180):

yup

Zeronknight (zeronknight):

What next XD

OpenStudy (legomyego180):

divide by 15.

Zeronknight (zeronknight):

16?

Zeronknight (zeronknight):

I think my teacher messed up or something, because it has no 16 in it lol

OpenStudy (legomyego180):

No I messed up. One sec.

OpenStudy (legomyego180):

\[\frac{ 12 }{ 15 } = \frac{ 20 }{ x }\] So you do 20 times 15 then divide by 12

OpenStudy (legomyego180):

So a is the answer, 25